OpenStudy (anonymous):

I have two problems to solve and I need help. 1. X^4-2x^2+64 2. X^4-52x^2+576

OpenStudy (anonymous):

the second problem use y = x^2 to make the equation y^2 -52y +576 = 0

OpenStudy (anonymous):

u should get y = 36 or y = 16 then sub y back into x^2 = y for example would be x^2 = 36 or x^2 = 16, so u should get x = 4 orx = -4 or x= 6 or x= -6